Grade 12Physical Chemistry

A non-volatile solute of 10gm is dissolved in 180 gm of water. If the relative lowering of vapour pressure is 0.005 then calculate the molecular weight of the solute.

Anirban Mukherjee
8 Years agoGrade 12

2 Answers

Arun
8 Years ago
Dear student
From Roult's law,
Relative lowering in vapour pressure = \Deltap = p° -p/p°
= n/N = w *M/m* W
w = 10 gm, W = 180 gm, m = ?, M = 18 gm
\Deltap = 0.005
Hence
0.005 = 10*18/180* m
m = 1/0.005 = 200
Regards
Arun (askIITians forum expert)
Shiva Saraswat
7 Years ago
10gm of non volatile solute is dissolved in 180 gm of water resulting in lowering of vapour pressure by 0.5℅.determine the boiling point of solution point of solution of kb of water is 0.52kkg/mole.